OpenStudy (lgbasallote):

\[\begin{array}{cc} \text{Find} \; \mathcal L \lbrace \phi (t) \rbrace \; \text{where} \; \phi (t) &= 1 \quad 0 < t < 2\\&=t \quad \quad \; \; t >2 \end{array}\]

OpenStudy (alexwee123):

uhh o.0 it looks hard is this like parametrics ? :o

OpenStudy (lgbasallote):

this is laplace transform haha lol

OpenStudy (anonymous):

Try the same here.. As you did earlier..

OpenStudy (anonymous):

\[\large \implies \int\limits_{0}^{2}(e^{-st})dt + \int\limits _{2}^{\infty} (t \cdot e^{-st}) dt\]

OpenStudy (lgbasallote):

uhh wait...i posted the wrong question...

OpenStudy (lgbasallote):

nevermind...i'll just solve this one

OpenStudy (anonymous):

You will be having problem in second integral I guess.. If no then Very Good for you..

OpenStudy (lgbasallote):

i got something like \[\huge \left[-\frac{e^{-st}}{s}\right|_0^2 + \left[-\frac{te^{-st}}{s} - \frac{e^{-st}}{s^2}\right|_2^\infty\] is that right?

OpenStudy (anonymous):

Sorry here it is t so you can solve it easily sorry.. But if it was t^{-1} then you will be having problems in evaluating this..

OpenStudy (anonymous):

Let me check.

OpenStudy (lgbasallote):

...i have no idea what you just said...

OpenStudy (lgbasallote):

@waterineyes still here?

OpenStudy (lgbasallote):

i got \[\huge \frac{e^{-2s} + 1}{s} + \frac{e^{-2s}}{s^2}\] can anyone check if im right?

OpenStudy (lgbasallote):

so wolfram can solve laplace o.O >.<

OpenStudy (unklerhaukus):

wolfram can do al lot of things http://www.wolframalpha.com/input/?i=blue+%2B+red

OpenStudy (lgbasallote):

but how did you know my answer is right?

OpenStudy (unklerhaukus):

i compared your piecewise function to the graph

OpenStudy (lgbasallote):

oh. well thanks

OpenStudy (anonymous):

My internet sucked that time... So I was unable to connect...

OpenStudy (lgbasallote):

maybe it was just OS...we never know